OpenStudy (anonymous):

Algebra 2 If f(x) = x2 + 2x + 3, what is the average rate of change of f(x) over the interval [-4, 6]? A. 51 B. 40 C. 31 D. 20 E. 4 is it E?

OpenStudy (freckles):

just evaluate \[\frac{f(6)-f(-4)}{6-(-4)}\]

OpenStudy (anonymous):

is it 20

OpenStudy (freckles):

what is f(6)=? and what is f(4)=?

OpenStudy (freckles):

oops f(-4)?

OpenStudy (freckles):

\[f(x)=x^2+2x+3 \\ f(6)=6^2+2(6)+3=? \\ f(-4)=(-4)^2+2(-4)+3=?\]

OpenStudy (anonymous):

uhh was it 51 and 4 ?

OpenStudy (freckles):

(-4)^2=(-4)(-4)=16 2(-4)=-8 16+(-8)+3 8+3 11 I get 11 for f(-4)

OpenStudy (freckles):

\[\frac{f(6)-f(-4)}{6-(-4)}=\frac{51-11}{10}\]

OpenStudy (anonymous):

so was I right then with the final answer being 4... I really suck at this...

OpenStudy (freckles):

yes 51-11=40 and 40/10=4

OpenStudy (anonymous):

Thank you

OpenStudy (freckles):

np